A block of metal is 5 cm by 7 cm by 9 cm and has a mass of 854 grams. Calculate the density of the block and give your best guess of what it is made of. Express the density in both g/cm3 and kg/m3.
1
Expert's answer
2019-04-24T09:56:48-0400

Let's first find the volume of the block:


"V = a \cdot b \cdot c = 5 cm \cdot 7 cm \cdot 9 cm = 315 cm^3."

Then, we can find the density of the block of metal from the formula:


"\rho = \dfrac{m}{V} = \dfrac{854 g}{315 cm^3} = 2.71 \dfrac{g}{cm^3}."

Let's, also, convert "g\/cm^3" to "kg\/m^3":


"\rho = 2.71 \dfrac{g}{cm^3} \cdot \dfrac{1 kg}{1000 g} \cdot \dfrac{(100 cm)^3}{1 m^3} = 2710 \dfrac{kg}{m^3}."

As we can see from the table [1], the metal is made of aluminum.

[1]. https://www.engineeringtoolbox.com/metal-alloys-densities-d_50.html

Answer:

"\rho = 2.71 \dfrac{g}{cm^3} = 2710 \dfrac{kg}{m^3}."